4. Inverse operations is a concept that needs to be

understood because it is often used in algebra to

solve variables.

We explain inverse operations as applied in numerical

statements. Inspect the following example of a

numerical equation, which is later translated into an

algebraic equation:

21 × 5 - 10 = 95

21 × 5 = 95 + 10 (Add 10 both sides of =.)

21 = (Divide by 5 both sides of =.)



Do the following:

Following our example to

the left, work through the

numerical equation below

and then translate it into an

algebraic equation, where you

solve for the unknown value:

35/5+ 13 = 20
1
Expert's answer
2021-06-01T07:26:38-0400

Answer:-

"\frac{3}{5}x+13=20"

"\frac{3}{5}x+13-13=20-13"

"\frac{3}{5}x=7"

"x=7\times\frac{5}{3}"

"x=\frac{35}{3}"

"x=11.67"